# Spatial-Mamba: Effective Visual State Space Models via Structure-aware State Fusion

### v1

@@ 4.1 Formulation of Spatial-Mamba
where $x_{t}$ is the original state variable, $h_{t}$ is the structure-aware state variable, $\Omega$ is the neighbor set, and $\rho_{k}(t)$ is the index of the $k$-th neighbor of position $t$. Fig. 2(b) illustrates the SSM flow in the proposed Spatial-Mamba. Compared with the original Mamba in Fig. 2(a), we can see that the original state variable $x_{t}$ is directly influenced by its previous state $x_{t-1}$, while the structure-aware state variable $h_{t}$ incorporates additional neighboring state variables $x_{\rho_{1}(t)},x_{\rho_{2}(t)},\dots,x_{\rho_{K}(t)}$ through a fusion mechanism, where $K=|\Omega|$ denotes the size of neighbor set. By considering both temporal and spatial information, the fused state variable $h_{t}$ gains a richer context, leading to improved adaptability and a more comprehensive understanding of the image.
@@
To gain an intuitive understanding of SASF, we visualize the original state variables and our proposed structure-aware state variables in Fig. 3. One can see that the original state $x_{t}$ (Fig. 3(b)) struggles to differentiate between the foreground and background. In contrast, the structure-aware state $h_{t}$ (Fig. 3(c)), which has been refined through a SASF process, effectively separates these regions. Moreover, the original state variable $x_{t}$ in Fig. 3(d) only shows horizontal attenuation along the scanning direction (gradually darkening from the brightest value in the upper left corner), while the fused state variable $h_{t}$ in Fig. 3(e) demonstrates decay along the horizontal, vertical and diagonal directions. This improvement stems from its ability to leverage spatial relationships within the image, leading to a more accurate and context-aware representation.
@@ 4.2 Network Architecture
The overall architecture of Spatial-Mamba is depicted in Fig. 4. It consists of four successive stages, resembling the structure of Swin-Transformer (Liu et al., 2021). We introduce three variants of Spatial-Mamba model at different scales: Spatial-Mamba-T (tiny), Spatial-Mamba-S (small), and Spatial-Mamba-B (base). The detailed configurations are provided in Appendix A. Specifically, an input image $I\in\mathbb{R}^{H\times W\times 3}$ is first processed by an overlapped stem layer to generate a 2D feature map with dimension of $\frac{H}{4}\times\frac{W}{4}\times C$. This feature map is then fed into four successive stages. Each stage comprises multiple Spatial-Mamba blocks, followed by a down-sampling layer with a factor of 2 (except for the last stage), resulting in hierarchical features. Finally, a head layer is employed to process these features to produce corresponding image representations for specific downstream tasks.
@@ 4.3 Connection with Original Mamba and Linear Attention
We analyze in-depth the similarities and disparities among linear attention (Katharopoulos et al., 2020), original Mamba (Gu & Dao, 2023) and our Spatial-Mamba, providing a better understanding of the working mechanism of our proposed method. Detailed derivations are provided in Appendix B.
@@
(a) Linear attention
(b) Mamba
(c) Spatial-Mamba
Figure 5: Visualization of matrices ${\bm{M}}$ for linear attention, Mamba and Spatial-Mamba.
Remarks. From the above analysis, we can see that all the three paradigms — linear attention, Mamba, and Spatial-Mamba — can be modeled within a unified matrix multiplication framework, specifically $y={\bm{M}}u$. The differences lie in the structure of ${\bm{M}}$. For both linear attention and Mamba, ${\bm{M}}$ takes the form of a lower triangular matrix, whereas for Spatial-Mamba, ${\bm{M}}$ is an adjacency matrix. Fig. 5 provides a visualization of these matrices. In linear attention, the positions of brighter values remain consistent along the vertical direction, which indicates that the SA mechanism puts its focus on a small set of image tokens. Mamba, on the other hand, shows a decaying pattern over time, which is attributed to the influence of its state transition matrix $\overline{{\bm{A}}}_{t}$. This dynamic transition allows Mamba to shift its focus among previous image tokens. Unlike linear attention and Mamba, our Spatial-Mamba considers the weighted summation of all states within a broader spatial neighborhood $\Omega$, allowing for a more comprehensive representation of spatial relationships.
@@ 5 Experimental Results
In this section, we conduct a series of experiments to compare Spatial-Mamba with leading benchmark models, including those based on Convolutional Neural Networks (CNNs) (Radosavovic et al., 2020; Liu et al., 2022; Yu & Wang, 2024), Vision Transformers (Dosovitskiy et al., 2020; Liu et al., 2021; Wang et al., 2022; Hassani et al., 2023), and recent visual SSMs (Nguyen et al., 2022; Zhu et al., 2024; Liu et al., 2024; Huang et al., 2024). Following previous works (Liu et al., 2021; 2024), we train three variants of Spatial-Mamba, namely Spatial-Mamba-T, Spatial-Mamba-S and Spatial-Mamba-B. The detailed configurations are provided in Appendix A. The performance evaluation is conducted on fundamental visual tasks, including image classification, object detection, and semantic segmentation.
@@ 5.1 Image Classification on ImageNet-1K
Settings. We first evaluate the representation learning capabilities of Spatial-Mamba in image classification on ImageNet-1K (Deng et al., 2009). We adopted the experimental configurations used in previous works (Liu et al., 2021; 2024), which are detailed in Appendix A. We compare our method with state-of-the-art approaches, including RegNetY (Radosavovic et al., 2020), ConvNeXt (Liu et al., 2022), ViT (Dosovitskiy et al., 2020), DeiT (Touvron et al., 2021), Swin (Liu et al., 2021), NAT (Hassani et al., 2023), S4ND (Nguyen et al., 2022), Vim (Zhu et al., 2024), VMamba (Liu et al., 2024), and LocalVMamba (Huang et al., 2024).
@@
Results. Tab. 1 presents a comprehensive comparison between Spatial-Mamba against state-of-the-art methods. Notably, Spatial-Mamba-T achieves a top-1 accuracy of 83.5%, outperforming the CNN-based ConvNext-T by 1.4% with similar amount of parameters and GFLOPs. Compared to Transformer-based methods, Spatial-Mamba-T exceeds Swin-T by 2.2% and NAT-T by 0.3%. In comparison with SSM-based methods, Spatial-Mamba-T outperforms VMamba-T by 1.0% and LocalVMamba-T by 0.8%. For other variants, Spatial-Mamba also shows advantages. Specifically, Spatial-Mamba-S and Spatial-Mamba-B achieve top-1 accuracies of 84.6% and 85.3%, respectively, surpassing NAT-S and NAT-B by margins of 1.6% and 1.0%, and VMamba-S and VMamba-B by 1.0% and 1.4%. While Spatial-Mamba-T is slightly slower than VMamba-T due to architectural differences, the Small and Base variants of Spatial-Mamba are faster than their VMamba counterparts. Moreover, both of them are significantly faster than CNN and Transformer-based methods.
@@ 5.2 Object Detection and Instance Segmentation on COCO
Settings. We evaluate Spatial-Mamba in object detection and instance segmentation tasks using COCO 2017 dataset (Lin et al., 2014) and MMDetection library (Chen et al., 2019). We adopt Mask R-CNN (He et al., 2017) as detector head, apply Spatial-Mamba-T/S/B pre-trained on ImageNet-1K as backbones. Following common practices (Liu et al., 2021; 2024), we fine-tune the pre-trained models for 12 epochs (1$\times$ schedule) and 36 epochs with multi-scale inputs (3$\times$ schedule). During training, AdamW optimizer is adopted with an initial learning rate of 0.0001 and a batch size of 16.
Results. Detailed results on COCO are reported in Tab. 2. It can be seen that all variants of Spatial-Mamba outperform their competitors under different schedules. For 1$\times$ schedule, Spatial-Mamba-T achieves a box mAP of 47.6 and a mask mAP of 42.9, surpassing Swin-T/VMamba-T by 4.9/0.3 in box mAP and 3.6/0.2 in mask mAP with fewer parameters and FLOPs, respectively. Similarly, Spatial-Mamba-S/B demonstrate superior performance to other methods under the same configuration. Furthermore, these trends of improved performance hold with the 3$\times$ multi-scale training schedule. Notably, Spatial-Mamba-S achieves the highest box mAP of 50.5 and mask mAP of 44.6, surpassing VMamba-S with a considerable gain of 0.6 and 0.4, respectively.
5.3 Semantic Segmentation on ADE20K
@@ 5.4 Ablation Studies
Neighbor set. First, adjustments to the neighbor set $\Omega$ reveal that increasing the size from a $3\times 3$ neighbourhood to $5\times 5$ results in a gradual improvement in accuracy, from 82.3% to 82.5%, albeit with a corresponding decrease in throughput. Furthermore, employing a broader dilated neighbor set with factors $d=1,3,5$ increases the accuracy to 82.7%, while reducing throughput to 1158. This suggests a trade-off between speed and larger neighbor set.
@@
In addition, we also provide some qualitative results in Appendix C, and a comparative analysis of the Effective Receptive Fields (ERF) (Ding et al., 2022) of various models is provided in Appendix D.
@@ 6 Conclusion
Appendix A Detailed Experiment Settings
Network architecture. The detailed architecture of Spatial-Mamba models are outlined in Tab. 5. Following the common four-stage hierarchical framework(Liu et al., 2021; Han et al., 2024), we constructed the Spatial-Mamba models by stacking proposed Spatial-Mamba blocks at each stage. Specifically, an input image with resolution of $224\times 224$ is firstly processed by a stem layer, which consists of convolutions, Batch Normalization (BN) layers and GELU activation function. The kernel size is $3\times 3$ with a stride of 2 at first and last convolutions and a stride of 1 for the others. Each stage contains multiple Spatial-Mamba blocks, following by a down-sampling layer except the last one. The down-sampling layer consists of a $3\times 3$ convolution with a stride of 2 and a Layer Normalization (LN) layer. Each block incorporates a structure-aware SSM layer and a Feed-Forward Network (FFN), both with residual connections. The structure-aware SSM contains a SASF branch with a 2D depth-wise convolution and a multiplicative gate branch with activation function, as illustrated in Fig. 4. The expand ratio of SSM is set to 2, doubling the number of channels. We modify the embedding dimension and number of blocks to build our Spatial-Mamba-T/S/B models.
Settings for ImageNet-1K classification. The Spatial-Mamba-T/S/B models are trained from scratch for 300 epochs using AdamW optimizer with betas set to (0.9, 0.999), momentum of 0.9 and batch size of 1024. The initial learning rate is set to 0.001 with a weight decay of 0.05 and a cosine annealing learning rate schedule is adopted with a warm-up of 20 epochs. We adopt the common data augmentation strategies consisted with previous work(Liu et al., 2021; 2024). Moreover, label smoothing (0.1), exponential moving average (EMA) and MESA(Du et al., 2022) are also applied. The drop path rate is set to 0.2 for Spatial-Mamba-T, 0.3 for Spatial-Mamba-S and 0.5 for Spatial-Mamba-B.
@@ Appendix A Detailed Experiment Settings
Appendix B Derivation of SSM Formulas
Recall the definition of Mamba in Eq. (1), by defining the initial state $x_{0}$ as zero, the state transition equation in the recursive form can be rewritten as follows:
@@ Appendix B Derivation of SSM Formulas
$=\overline{{\bm{A}}}_{t}\dots\overline{{\bm{A}}}_{1}x_{0}+\overline{{\bm{A}}}_{t}\dots\overline{{\bm{A}}}_{2}\overline{{\bm{B}}}_{1}u_{1}+\cdots+\overline{{\bm{A}}}_{t}\overline{{\bm{A}}}_{t-1}\overline{{\bm{B}}}_{t-2}u_{t-2}+\overline{{\bm{A}}}_{t}\overline{{\bm{B}}}_{t-1}u_{t-1}+\overline{{\bm{B}}}_{t}u_{t}$
$=\sum_{s\leq t}\overline{{\bm{A}}}_{s:t}^{\times}\ \overline{{\bm{B}}}_{s}u_{s},\quad\text{where }\overline{{\bm{A}}}_{s:t}^{\times}:=\begin{cases}\overline{{\bm{A}}}_{t}\overline{{\bm{A}}}_{t-1}\dots\overline{{\bm{A}}}_{s+1}&s<t\\ 1&s=t\end{cases}.$
@@
By omitting term ${\bm{D}}_{t}u_{t}$ for simplicity, and multiplying by ${\bm{C}}_{t}$ according to the observation equation, we derive the final output $y_{t}=\sum_{s\leq t}{\bm{C}}_{t}\ \overline{{\bm{A}}}_{s:t}^{\times}\ \overline{{\bm{B}}}_{s}u_{s}$. Defining the input vector $u=\left[u_{1},\cdots,u_{L}\right]^{T}\in\mathbb{R}^{L}$ with sequences length $L$ and corresponding output vector $y\in\mathbb{R}^{L}$, then the above calculation can be written with in matrix multiplication framework, i.e. $y={\bm{M}}u$, where ${\bm{M}}$ is a structured lower triangular matrix and ${\bm{M}}_{ij}={\bm{C}}_{i}\overline{{\bm{A}}}_{j:i}^{\times}\overline{{\bm{B}}}_{j}$.
@@
Similarly, we can also represent Spatial-Mamba in the same matrix transformation form. Based on Eq. (2) and Eq. (7), the SASF equation can be rewritten as $h_{t}=\sum_{k\in\Omega}\sum_{s\leq\rho_{k}(t)}\alpha_{k}\overline{{\bm{A}}}_{s:\rho_{k}(t)}^{\times}\overline{{\bm{B}}}_{s}u_{s}$. Then by multiplying ${\bm{C}}_{t}$ with out term ${\bm{D}}_{t}u_{t}$, we derive the final output of Spatial-Mamba, i.e. $y_{t}=\sum_{k\in\Omega}\sum_{s\leq\rho_{k}(t)}\alpha_{k}{\bm{C}}_{t}\overline{{\bm{A}}}_{s:\rho_{k}(t)}^{\times}\overline{{\bm{B}}}_{s}u_{s}$. This can also be concisely represented as a matrix multiplication form $y={\bm{M}}u$, where ${\bm{M}}$ is a structured adjacency matrix and ${\bm{M}}_{ij}=\sum_{k}{\bm{C}}_{i}\overline{{\bm{A}}}_{j:\rho_{k}(i)}^{\times}\overline{{\bm{B}}}_{j}$.
Appendix C Effective Receptive Field (ERF)
We compare the Effective Receptive Field (ERF) of the center pixel on popular backbone networks before and after training, as shown in Fig. 6. Our Spatial-Mamba-T and Deit-S, Vim-S and VMamba-T all show a global ERF. In addition, after training, both Vim-S and VMamba-T exhibit noticeable accumulation contributions along either horizontal or vertical directions, which is attributed to their multi-directional fusion mechanisms. In contrast, our unidirectional Spatial-Mamba-T effectively eliminates this phenomenon.
Appendix D More Qualitative Results

### v2

@@ 4.1 Formulation of Spatial-Mamba
where $x_{t}$ is the original state variable, $h_{t}$ is the structure-aware state variable, $\Omega$ is the neighbor set, $\alpha_{k}$ is a learnable weight, and $\rho_{k}(t)$ is the index of the $k$-th neighbor of position $t$. Fig. 2(b) illustrates the SSM flow in the proposed Spatial-Mamba. Compared with the original Mamba in Fig. 2(a), we can see that the original state variable $x_{t}$ is directly influenced by its previous state $x_{t-1}$, while the structure-aware state variable $h_{t}$ incorporates additional neighboring state variables $x_{\rho_{1}(t)},x_{\rho_{2}(t)},\dots,x_{\rho_{K}(t)}$ through a fusion mechanism, where $K=|\Omega|$ denotes the size of neighbor set. By considering both the global long-range and the local spatial information, the fused state variable $h_{t}$ gains a richer context, leading to improved adaptability and a more comprehensive understanding of the image.
@@
To gain an intuitive understanding of SASF, we visualize the original state variables and our proposed structure-aware state variables in Fig. 3. Additional visualizations are provided in Appendix A. One can see that the original state $x_{t}$ (Fig. 3(b)) struggles to differentiate the foreground from background. In contrast, the structure-aware state $h_{t}$ (Fig. 3(c)), which has been refined through a SASF process, effectively separates these regions. Moreover, the original state variable $x_{t}$ in Fig. 3(d) only shows horizontal attenuation along the scanning direction (gradually darkening from the brightest value in the upper left corner), while the fused state variable $h_{t}$ in Fig. 3(e) demonstrates decay along the horizontal, vertical and diagonal directions. This improvement stems from its ability to leverage spatial relationships within the image, leading to a more accurate and context-aware representation.
@@ 4.2 Network Architecture
The overall architecture of Spatial-Mamba is depicted in Fig. 4. It consists of four successive stages, resembling the structure of Swin-Transformer (Liu et al., 2021). We introduce three variants of Spatial-Mamba model at different scales: Spatial-Mamba-T (tiny), Spatial-Mamba-S (small), and Spatial-Mamba-B (base). The detailed configurations are provided in Appendix B. Specifically, an input image $I\in\mathbb{R}^{H\times W\times 3}$ is first processed by an overlapped stem layer to generate a 2D feature map with dimension of $\frac{H}{4}\times\frac{W}{4}\times C$. This feature map is then fed into four successive stages. Each stage comprises multiple Spatial-Mamba blocks, followed by a down-sampling layer with a factor of 2 (except for the last stage), resulting in hierarchical features. Finally, a head layer is employed to process these features to produce corresponding image representations for specific downstream tasks.
@@ 4.3 Connection with Original Mamba and Linear Attention
We analyze in-depth the similarities and disparities among linear attention (Katharopoulos et al., 2020), original Mamba (Gu & Dao, 2023) and our Spatial-Mamba, providing a better understanding of the working mechanism of our proposed method. Detailed derivations are provided in Appendix C.
@@
(a) Input
(b) Linear attention
(c) Mamba
(d) Spatial-Mamba
Figure 5: Visualizations of matrices ${\bm{M}}$ and the corresponding activation maps for linear attention, Mamba and Spatial-Mamba. The red arrows indicate specific rows in matrices ${\bm{M}}$, along with the corresponding image patches (marked with a red star).
Remarks. From the above analysis, we can see that all the three paradigms — linear attention, Mamba, and Spatial-Mamba — can be modeled within a unified matrix multiplication framework, specifically $y={\bm{M}}u$. The differences lie in the structure of ${\bm{M}}$. For both linear attention and Mamba, ${\bm{M}}$ takes the form of a lower triangular matrix, whereas for Spatial-Mamba, ${\bm{M}}$ is an adjacency matrix. Fig. 5 provides a visualization of these matrices. In linear attention, the positions of brighter values remain consistent along the vertical direction, which indicates that the SA mechanism puts its focus on a small set of image tokens. Mamba, on the other hand, shows a decaying pattern over time, which is attributed to the influence of its state transition matrix $\overline{{\bm{A}}}_{t}$. This dynamic transition allows Mamba to shift its focus among previous image tokens. Unlike linear attention and Mamba, our Spatial-Mamba considers the weighted summation of all states within a broader spatial neighborhood $\Omega$, allowing for a more comprehensive representation of spatial relationships. The activation maps on the right side of Fig. 5 further demonstrate that linear attention focuses on a limited region, while Mamba captures a broader region due to its long-range context modeling. Our Spatial-Mamba not only largely extends the range of context modeling but also enables spatial structure modeling, effectively identifying relevant regions even when they are distant from each other.
@@ 5 Experimental Results
In this section, we conduct a series of experiments to compare Spatial-Mamba with leading benchmark models, including those based on Convolutional Neural Networks (CNNs) (Radosavovic et al., 2020; Liu et al., 2022; Yu & Wang, 2024), Vision Transformers (Dosovitskiy et al., 2020; Liu et al., 2021; Wang et al., 2022; Hassani et al., 2023), and recent visual SSMs (Nguyen et al., 2022; Zhu et al., 2024; Liu et al., 2024; Huang et al., 2024). Following previous works (Liu et al., 2021; 2024), we train three variants of Spatial-Mamba, namely Spatial-Mamba-T, Spatial-Mamba-S and Spatial-Mamba-B. The detailed configurations are provided in Appendix B. The performance evaluation is conducted on fundamental visual tasks, including image classification, object detection, and semantic segmentation.
@@ 5.1 Image Classification on ImageNet-1K
Settings. We first evaluate the representation learning capabilities of Spatial-Mamba in image classification on ImageNet-1K (Deng et al., 2009). We adopted the experimental configurations used in previous works (Liu et al., 2021; 2024), which are detailed in Appendix B. We compare our method with state-of-the-art approaches, including RegNetY (Radosavovic et al., 2020), ConvNeXt (Liu et al., 2022), ViT (Dosovitskiy et al., 2020), DeiT (Touvron et al., 2021), Swin (Liu et al., 2021), NAT (Hassani et al., 2023), S4ND (Nguyen et al., 2022), Vim (Zhu et al., 2024), VMamba (Liu et al., 2024), and LocalVMamba (Huang et al., 2024).
@@
Results. Tab. 1 presents a comprehensive comparison between Spatial-Mamba against state-of-the-art methods. Notably, Spatial-Mamba-T achieves a top-1 accuracy of 83.5%, outperforming the CNN-based ConvNext-T by 1.4% with similar amount of parameters and FLOPs. Compared to Transformer-based methods, Spatial-Mamba-T exceeds Swin-T by 2.2% and NAT-T by 0.3%. In comparison with SSM-based methods, Spatial-Mamba-T outperforms VMamba-T by 1.0% and LocalVMamba-T by 0.8%. For other variants, Spatial-Mamba also shows advantages. Specifically, Spatial-Mamba-S and Spatial-Mamba-B achieve top-1 accuracies of 84.6% and 85.3%, respectively, surpassing NAT-S and NAT-B by margins of 1.6% and 1.0%, and VMamba-S and VMamba-B by 1.0% and 1.4%. While Spatial-Mamba-T is slightly slower than VMamba-T due to architectural differences, the Small and Base variants of Spatial-Mamba are faster than their VMamba counterparts. Moreover, both of them are significantly faster than CNN and Transformer-based methods.
@@ 5.2 Object Detection and Instance Segmentation on COCO
Settings. We evaluate Spatial-Mamba in object detection and instance segmentation tasks using COCO 2017 dataset (Lin et al., 2014) and MMDetection library (Chen et al., 2019). We adopt Mask (He et al., 2017) and Cascade Mask R-CNN (Cai & Vasconcelos, 2018) as detector heads, apply Spatial-Mamba-T/S/B pre-trained on ImageNet-1K as backbones. Following common practices (Liu et al., 2021; 2024), we fine-tune the pre-trained models for 12 epochs (1$\times$ schedule) and 36 epochs with multi-scale inputs (3$\times$ schedule). During training, AdamW optimizer is adopted with an initial learning rate of 0.0001 and a batch size of 16.
Results. The results on COCO with Mask R-CNN are reported in Tab. 2, and the results with Cascade Mask R-CNN are provided in Appendix D. It can be seen that all variants of Spatial-Mamba outperform their competitors under different schedules. For 1$\times$ schedule, Spatial-Mamba-T achieves a box mAP of 47.6 and a mask mAP of 42.9, surpassing Swin-T/VMamba-T by 4.9/0.3 in box mAP and 3.6/0.2 in mask mAP with fewer parameters and FLOPs, respectively. Similarly, Spatial-Mamba-S/B demonstrate superior performance to other methods under the same configuration. Furthermore, these trends of improved performance hold with the 3$\times$ multi-scale training schedule. Notably, Spatial-Mamba-S achieves the highest box mAP of 50.5 and mask mAP of 44.6, surpassing VMamba-S with a considerable gain of 0.6 and 0.4, respectively.
@@ 5.4 Ablation Studies
Neighbor set. First, adjustments to the neighbor set $\Omega$ reveal that increasing the size from a $3\times 3$ neighborhood to $5\times 5$ results in a gradual improvement in accuracy, from 82.3% to 82.5%, albeit with a corresponding decrease in throughput. Furthermore, employing a broader dilated neighbor set with factors $d=1,3,5$ increases the accuracy to 82.7%, while reducing throughput to 1158. This suggests a trade-off between speed and larger neighbor set.
@@
In addition, we provide some qualitative results in Appendix E, an in-depth discussion about SASF in Appendix F, and a comparative analysis of the Effective Receptive Fields (ERF) (Ding et al., 2022) of various models is provided in Appendix G.
@@ 6 Conclusion
Appendix to “Spatial-Mamba: Effective Visual State Space
Models via Structure-aware State Fusion”
In this appendix, we provide the following materials:
  • A
    More visual results for SASF (referring to Sec. 4.1 in the main paper);
  • B
    Architecture details of Spatial-Mamba and experimental settings (referring to Sec. 4.2 and Sec. 5.1 in the main paper);
  • C
    Derivations of the SSM formulas of Mamba and Spatial-Mamba (referring to Sec. 4.3 in the main paper);
  • D
    Results of Cascade Mask R-CNN detector head (referring to Sec. 5.2 in the main paper);
  • E
    Visual results and comparisons for detection and segmentation tasks (referring to Sec. 5.2 and Sec. 5.3 in the main paper);
  • F
    Further discussions about SASF (referring to Sec. 5.4 in the main paper);
  • G
    Comparisons of effective receptive field (referring to Sec. 5.4 in the main paper).
Appendix A More Visual Results
More visual comparisons between the original and fused structure-aware state variables are shown in Fig. 6. Due to the spatial modeling capability of SASF, the fused state variables demonstrate more accurate context information and superior structural perception across different scenarios.
(a) Input
(b) Original state
(c) Fused state
(d) Input
(e) Original state
(f) Fused state
Figure 6: More visualizations of state variables before and after applying the SASF equation.
Appendix B Model Architectures and Experiment Settings in Classification
@@
Network architecture. The detailed architectures of Spatial-Mamba models are outlined in Tab. 5. Following the common four-stage hierarchical framework (Liu et al., 2021; Han et al., 2024), we construct the Spatial-Mamba models by stacking our proposed Spatial-Mamba blocks at each stage. Specifically, an input image with resolution of $224\times 224$ is firstly processed by a stem layer, which consists of Convolution (Conv), Batch Normalization (BN) and GELU activation function. The kernel size is $3\times 3$ with a stride of 2 at the first and last convolution layers, and a stride of 1 for other layers. Each stage contains multiple Spatial-Mamba blocks, followed by a down-sampling layer except for the last block. The down-sampling layer consists of a $3\times 3$ convolution with a stride of 2 and a Layer Normalization (LN) layer. Each block incorporates a structure-aware SSM layer and a Feed-Forward Network (FFN), both with residual connections. The structure-aware SSM contains a SASF branch with a 2D depth-wise convolution and a multiplicative gate branch with activation function, as illustrated in Sec. 4.2 of the main paper. The expand ratio of SSM is set to 2, doubling the number of channels. The SSM state dimension is set to 1 for better performance and efficiency. We modify the embedding dimension and number of blocks to build our Spatial-Mamba-T/S/B models.
Settings for ImageNet-1K classification. The Spatial-Mamba-T/S/B models are trained from scratch for 300 epochs using AdamW optimizer with betas set to (0.9, 0.999), momentum set to 0.9, and batch size set to 1024. The initial learning rate is set to 0.001 with a weight decay of 0.05. A cosine annealing learning rate schedule is adopted with a warm-up of 20 epochs. We adopt the common data augmentation strategies as in previous works (Liu et al., 2021; 2024). Moreover, label smoothing (0.1), exponential moving average (EMA) and MESA (Du et al., 2022) are also applied. The drop path rate is set to 0.2 for Spatial-Mamba-T, 0.3 for Spatial-Mamba-S and 0.5 for Spatial-Mamba-B.
Implementation details. The Spatial-Mamba models employ a hardware-aware selective scan algorithm adapted from the original Mamba framework, with modifications to the CUDA kernels for decoupling state transition and observation equations. The SASF module of Spatial-Mamba is implemented by the general matrix multiplication (GEMM) with optimized CUDA kernels.
Appendix C Derivation of SSM Formulas
Based on the Mamba formulation provided in Sec. 3 of the main paper, the state transition equation in the recursive form can be rewritten as follows:
@@ Appendix C Derivation of SSM Formulas
$=\Pi_{i=1}^{t}\overline{{\bm{A}}}_{i}x_{0}+\Pi_{i=2}^{t}\overline{{\bm{A}}}_{i}\overline{{\bm{B}}}_{1}u_{1}+\cdots+\Pi_{i=t-1}^{t}\overline{{\bm{A}}}_{i}\overline{{\bm{B}}}_{t-2}u_{t-2}+\Pi_{i=t}^{t}\overline{{\bm{A}}}_{i}\overline{{\bm{B}}}_{t-1}u_{t-1}+\overline{{\bm{B}}}_{t}u_{t}$
By defining the initial state as zero, $x_{0}=0$, Eq. (7) can be expressed as:
$x_{t}$ $=\sum_{s\leq t}\overline{{\bm{A}}}_{s:t}^{\times}\ \overline{{\bm{B}}}_{s}u_{s},\quad\text{where }\overline{{\bm{A}}}_{s:t}^{\times}:=\begin{cases}\Pi_{i=s+1}^{t}\overline{{\bm{A}}}_{i},&s<t\\ 1,&s=t\end{cases}.$ (8)
@@
We omit the term ${\bm{D}}_{t}u_{t}$ for simplicity. According to the observation equation, we can derive the final output $y_{t}=\sum_{s\leq t}{\bm{C}}_{t}\ \overline{{\bm{A}}}_{s:t}^{\times}\ \overline{{\bm{B}}}_{s}u_{s}$. Suppose the input vector is $u=\left[u_{1},\cdots,u_{L}\right]^{T}\in\mathbb{R}^{L}$ with length $L$, the corresponding output vector is $y\in\mathbb{R}^{L}$. Then the above calculation can be written in matrix multiplication form, i.e., $y={\bm{M}}u$, where ${\bm{M}}$ is a structured lower triangular matrix and ${\bm{M}}_{ij}={\bm{C}}_{i}\overline{{\bm{A}}}_{j:i}^{\times}\overline{{\bm{B}}}_{j}$.
@@
Similarly, we can represent Spatial-Mamba in the same matrix transformation form. Based on the definition of Spatial-Mamba in Sec. 4.1 and Eq. (8), the SASF equation can be rewritten as $h_{t}=\sum_{k\in\Omega}\sum_{s\leq\rho_{k}(t)}\alpha_{k}\overline{{\bm{A}}}_{s:\rho_{k}(t)}^{\times}\overline{{\bm{B}}}_{s}u_{s}$. By multiplying ${\bm{C}}_{t}$, we can derive the final output of Spatial-Mamba as $y_{t}=\sum_{k\in\Omega}\sum_{s\leq\rho_{k}(t)}\alpha_{k}{\bm{C}}_{t}\overline{{\bm{A}}}_{s:\rho_{k}(t)}^{\times}\overline{{\bm{B}}}_{s}u_{s}$. It can also be concisely represented as a matrix multiplication form $y={\bm{M}}u$, where ${\bm{M}}$ is a structured adjacency matrix and ${\bm{M}}_{ij}=\sum_{k}{\bm{C}}_{i}\overline{{\bm{A}}}_{j:\rho_{k}(i)}^{\times}\overline{{\bm{B}}}_{j}$.
Appendix D Results of Cascade Mask R-CNN Detector Head
Detailed results of object detection and instance segmentation on the COCO dataset with Cascade Mask R-CNN framework are reported in Tab. 6. We can see that Spatial-Mamba-T achieves a box mAP of 52.1 and a mask mAP of 44.9, surpassing Swin-T/NAT-T by 1.7/0.7 in box mAP and 1.2/0.4 in mask mAP with fewer parameters and FLOPs, respectively. Similarly, Spatial-Mamba-S demonstrates superior performance under the same configuration.
Table 6: Results of object detection and instance segmentation on the COCO dataset using Cascade Mask R-CNN (Cai & Vasconcelos, 2018) under 3$\times$ schedule. FLOPs are calculated with input resolution of 1280 × 800.
[…]
Swin-T 50.4 69.2 54.7 43.7 66.6 47.3 86M 745G
ConvNeXt-T 50.4 69.1 54.8 43.7 66.5 47.3 86M 741G
NAT-T 51.4 70.0 55.9 44.5 67.6 47.9 85M 737G
Spatial-Mamba-T 52.1 71.0 56.5 44.9 68.3 48.7 84M 740G
Swin-S 51.9 70.7 56.3 45.0 68.2 48.8 107M 838G
ConvNeXt-S 51.9 70.8 56.5 45.0 68.4 49.1 108M 827G
NAT-S 52.0 70.4 56.3 44.9 68.1 48.6 108M 809G
Spatial-Mamba-S 53.3 71.9 57.9 45.8 69.4 49.7 101M 794G
Appendix E Qualitative Results